# Differentiable master equation solver for quantum device characterisation

### v1

@@ Appendix A Differentiable Solver Test Case
Figure 10: The total time (a) and the time per output (b) for computations on a range of batch sizes using the differentiable master equation for the single quantum dot model described in Eq. (A28). Each point is an average of 100 evaluation times and the time per output is simply the total time divided by the batch size. Output refers to the time taken to compute the output of a batch of input parameters. To compute a gradient by automatic differentiation, a variable must be tracked during the output computation, and once the output is available the gradient can be evaluated. Both plots share the same legend. The black lines corresponding to QuTiP execution times are extrapolated from the time taken for a single computation.
@@
Appendix B Algorithm Overview
An overview of our approach for dealing with a mixture of differentiable and non-differentiable parameters is shown in Figure 11. We show the order of operation used for the case studies discussed in the main text. After an exploratory log-spaced grid search, the optimisation of non-differentiable parameters uses a Nelder-Mead optimiser to minimise the loss from optimisation of the differentiable parameters using the given non-differentiable parameters. The final step is to perform a longer optimisation of differentiable parameter using the optimised non-differentiable parameters.
Figure 11: A schematic of our approach for optimising differentiable and non-differentiable parameters when fitting to data. The example shown is purely illustrative and does not contain real data. Here, we denote non-differentiable parameters as $\bm{\theta}^{\mathrm{v}}_{\mathrm{nd}}$ and differentiable parameters as $\bm{\theta}^{\mathrm{v}}_{\mathrm{diff}}$. Sets of optimised parameters are indicated with an asterisk. Optimisation of differentiable parameters is indicated by the schematic introduced in Figure 1 (a).
Appendix C Double Quantum Dot Parameter Estimation
@@
Table 2 displays the parameter estimation results for the example shown in Figure 4. The bias is set to $0.241\ \mathrm{mV}$, and the remaining parameters are set to representative fixed values of $c_{s}=3000\ \mathrm{ms}^{-1}$, $a=20\ \mathrm{nm}$, and $d=100\ \mathrm{nm}$. The performance of this example is representative of our approach’s parameter estimation for this model.
@@
Appendix D Spectral Densities
@@
$\lambda_{\mathbf{k}}=M_{\mathbf{k}}\big{(}1-e^{i\bm{\kappa}\cdot\bm{\rho}}\big{)}|\tilde{\psi}(\bm{\kappa},k_{z})|^{2}.$ (D37)
@@
$F(\omega)=\sum_{\mathbf{k}}\delta(\omega-ck)|\tilde{\psi}(\kappa,k_{z})|^{4}|1-e^{i\kappa\cdot\mathbf{d}}|^{2}.$ (D38)
@@
$\begin{split}F(\omega)=\frac{v\omega^{2}}{2\pi^{2}c_{s}^{3}}\int_{0}^{\pi}&|\tilde{\psi}\Big{(}\frac{\omega}{c_{s}}\sin{\theta},\frac{\omega}{c_{s}}\cos{\theta}\Big{)}|^{4}\times\\ &[1-J_{0}\Big{(}\frac{\omega d}{c_{s}}\sin{\theta}\Big{)}]\sin{\theta}d\theta,\end{split}$ (D39)
@@
$\begin{split}F(\omega)=\frac{v\omega^{2}}{\pi^{2}c_{s}^{3}}\int_{0}^{1}&\Big{|}\tilde{\psi}\Big{(}\frac{\omega}{c_{s}}\sqrt{1-x^{2}},\frac{\omega}{c_{s}}x\Big{)}\Big{|}^{4}\times\\ &\Big{[}1-J_{0}\Big{(}\frac{\omega d}{c_{s}}\sqrt{1-x^{2}}\Big{)}\Big{]}dx.\end{split}$ (D40)
Assuming a sharply peaked electron density simplifies the integral as $|\tilde{\psi}(\bm{\kappa},k_{z})|^{4}\approx 1$. In the context of a double quantum dot system this is a reasonable assumption as the dots are usually confined to a narrow region of space in all directions. The finite extent of the dot wavefunctions, assumed to be Gaussian $\psi(\mathbf{r})\propto e^{-r^{2}/2a^{2}}$, introduces a Gaussian decay from the Fourier transform of the dot wavefunction in D37.
@@
$F(\omega)=\frac{v}{\pi^{2}cd^{2}}\bigg{(}\frac{\omega}{\omega_{d}}\bigg{)}^{2}\bigg{(}1-\operatorname{sinc}{\frac{\omega}{\omega_{d}}}\bigg{)}e^{-\omega^{2}/2\omega_{a}^{2}},$ (D41)
@@
$M_{\mathbf{k}}=\bigg{(}\frac{1}{2M\omega_{\mathbf{k}}}\bigg{)}^{-1/2}(D|\mathbf{k}|+iP),$ (D42)

### v2

@@ Appendix A Differentiable Solver Test Case
Figure 10: Batch size scaling of direct computation of the steady state. The total time (a) and the time per output (b) for computations on a range of batch sizes using the differentiable master equation for the single quantum dot model described in Eq. (A28). Each point is an average of 100 evaluation times and the time per output is simply the total time divided by the batch size. Output refers to the time taken to compute the output of a batch of input parameters. To compute a gradient by automatic differentiation, a variable must be tracked during the output computation, and once the output is available the gradient can be evaluated. Both plots share the same legend. The black lines corresponding to QuTiP execution times are extrapolated from the time taken for a single computation.
@@
We also consider how the computation of the steady state scales with the dimension of the Hilbert space. This scaling is shown in Figure 11, where we find that our solver scales with Hilbert space size, $\mathcal{H}_{\mathrm{dim}}$, as $\sim\!\mathcal{H}_{\mathrm{dim}}^{4}$. This is as expected from the system of equations being solved in $\dot{\rho}(t)=0$ growing as $\mathcal{H}_{\mathrm{dim}}^{2}$. Given the time-scales shown in Figure 11, we expect our characterisation approach to be effective in practical times for systems of up to 6 dimensions, however larger systems could be considered if characterisation time is not a constraint. For the quantum dot systems considered in our case studies, a 6-dimensional system would equate to a chain of 5 quantum dots in the single excitation manifold.
Figure 11: Hilbert space scaling of the direct computation of the steady state. (a) Demonstrates the same outputs as in Figure 10 and how their evaluation time scales with Hilbert space size with a batch size of 100. (b) Displays the output time scaling with batch size and Hilbert space size. Each point in the average of 100 repetitions in (a) and 10 repetitions in (b). The Hilbert space dimension is increased by padding the basis vectors defining the single dot system described in Eq. (A28).
As our results consider steady state solutions, we compare the speed of finding the steady state using direct computation, as shown in Figure 10, with a long-time integration to the steady state using an ODE solver. As discussed in the main text, we implement a $4^{\mathrm{th}}$ order Runge-Kutta (RK4) method in TensorFlow for time dynamics for which the scaling with batch size is shown in Figure 12. It is clear from the comparison of batch size scaling in Figures 10 and 12 that direct computation of the steady state is much faster and more memory efficient than using an ODE solver. More advanced ODE solvers may find some improvement over our implementation, however due to the necessarily larger number of operations in evolving an ODE it is unlikely that they would achieve equivalent performance to direct computation of the steady state.
Figure 12: Batch size scaling of computing the steady state using an ODE solver. The total time (a) and the time per output (b) for computations on a range of batch sizes using the differentiable master equation for the single quantum dot model described in Eq. (A28) using an RK4 ODE solver. The ODE is solved for an initial state of $|1\rangle$ for 10 ns with 1000 time steps. Each point is an average of 100 evaluation times and the time per output is simply the total time divided by the batch size. Output refers to the time taken to compute the output of a batch of input parameters. Both plots share the same legend. The resource intensive nature of this computation leads to unexpected scaling beyond a batch size of $\sim\!10^{2}$, and the last two points in each series corresponding to gradient tracking and associated output are absent as the CPU resources were exhausted for these batch sizes.
Appendix B Algorithm Scaling
We display the scaling of each constituent part of our algorithm in Figure 13, namely the grid search over variables and final gradient descent optimisation. The grid search and gradient descent each take times of a similar magnitude, however the the grid search scales less favourably as can be expected from a simple grid search. Our algorithm has several hyper-parameters which can be tuned depending on requirements for speed and accuracy. The most relevant parameters for scaling are the number of points in the grid search over each variable, and the number of gradient descent steps in both the grid search and final gradient descent optimisation. The overall algorithm is then subject to the scaling of the Nelder-Mead optimiser. We also note that Bayesian optimisation is a valid alternative to a Nelder-Mead optimiser. We selected Nelder-Mead over Bayesian optimisation as the more reliable approach for our application. If one performs sequential estimates using Bayesian updates to priors using posteriors, the approach would scale linearly with the number of updates.
Figure 13: Time taken to perform (a) gradient descent optimisation of variables and (b) the grid search over variables for the model presented in Section VI. Each point uses the same algorithm hyper-parameters as for the results in Section VI and is the average of 10 repetitions. The relevant hyper-parameters are: in (a), 2500 gradient descent steps, and in (b), 7 points in the grid search for each variable and 400 gradient descent steps to compute the loss.
Gradient-free optimisation protocols, such as Nelder-Mead and Bayesian optimisation, do not scale well with number of parameters which may limit aspects of our approach to more complex systems. As discussed in Appendix D, we anticipate that breaking more complex systems into smaller characterisation steps with fewer parameters would circumvent this poor scaling.
Appendix C Errors and Posterior Filtering
The mean and standard deviation values for the distributions displayed in Figure 3 relating to the first case study are shown in Table 2. This table highlights that our parameter estimates are not strongly biased and have good consistency. The posterior mean errors have a lower standard deviation than optimal value errors, particularly in the case of temperature.
[…]
Parameter Optimal Posterior Mean
$\mu_{l}$ $-0.01\pm 0.64$
$\mu_{r}$ $-0.00\pm 0.08$
$\delta$ $-0.72\pm 3.59$
$\Gamma_{\mathrm{L}}$ $-0.60\pm 18.83$ $-0.87\pm 16.59$
$\Gamma_{\mathrm{R}}$ $-1.61\pm 12.62$ $-1.76\pm 12.24$
$T$ $0.96\pm 8.28$ $-0.09\pm 5.40$
Table 2: Summary statistics (mean and standard deviation) for the distributions displayed in Figure 3. As the standard deviation is sensitive to outliers, we remove values greater than 3 standard deviations away from the mean before recomputing the mean and standard deviation presented here. The number of outliers identified is $1\!-\!3\%$ for each distribution.
While the percentage error distributions for $\Gamma_{L}$ and $\Gamma_{R}$ have standard deviations which are not unreasonable for parameters which can vary by orders of magnitude in an experimental setting, the precision can be improved. To achieve this improvement, we capitalise on the extra information contained in the posterior distributions to reject samples with high parameter uncertainty. For each parameter posterior distribution, $P(\theta|D)$, we use the mean $\mu^{\theta}_{\mathrm{post}}$ and standard deviation $\sigma^{\theta}_{\mathrm{post}}$ to accept a result if all parameters fulfill the condition $|\sigma^{\theta}_{\mathrm{post}}/\mu^{\theta}_{\mathrm{post}}|\leq t$, where $t$ is a threshold value. By choosing $t\!=\!0.05$, the error distributions are significantly narrowed such that $\Gamma_{\mathrm{L}}:-0.81\pm 3.72\%$, $\Gamma_{\mathrm{R}}:1.58\pm 6.05\%$, and $T:0.17\pm 2.68\%$, while accepting $59\%$ of results. This is a useful approach when estimation accuracy and precision are prioritised, similar to threshold choices in classifiers to optimise true-positive rates craig2024bridging .
Figure 14: (a) The distribution of percentage errors over a test dataset using posterior mean estimates as a function of threshold value $t$ used for filtering results. For each parameter, the dashed line is the mean and the shaded area represents the mean plus/minus the standard deviation. (b) The percentage of accepted results using the filtering condition $|\sigma^{\theta}_{\mathrm{post}}/\mu^{\theta}_{\mathrm{post}}|\leq t$ for all differentiable parameters.
Appendix D Double Quantum Dot Parameter Estimation
@@
Table 3 displays the parameter estimation results for the example shown in Figure 4. The bias is set to $0.241\ \mathrm{mV}$, and the remaining parameters are set to representative fixed values of $c_{s}=3000\ \mathrm{ms}^{-1}$, $a=20\ \mathrm{nm}$, and $d=100\ \mathrm{nm}$. The performance of this example is representative of our approach’s parameter estimation for this model. Our approach is successful in identifying the values of $\epsilon_{0}$ and $\epsilon_{\mathrm{bias}}$ in this example, and the percentage error in these values across the case study test set is $3.48\%$ for $\epsilon_{0}$ and $-0.55\%$ for $\epsilon_{\mathrm{bias}}$. Identifying these points accurately and automatically is of benefit to experimental applications where the bias window must be known for pulsing qubits gorman2005charge ; petersson2010quantum ; hanson2007spins .
@@
Appendix E Spectral Densities
@@
$\lambda_{\mathbf{k}}=M_{\mathbf{k}}\big{(}1-e^{i\bm{\kappa}\cdot\bm{\rho}}\big{)}|\tilde{\psi}(\bm{\kappa},k_{z})|^{2}.$ (E37)
@@
$F(\omega)=\sum_{\mathbf{k}}\delta(\omega-ck)|\tilde{\psi}(\kappa,k_{z})|^{4}|1-e^{i\kappa\cdot\mathbf{d}}|^{2}.$ (E38)
@@
$\begin{split}F(\omega)=\frac{v\omega^{2}}{2\pi^{2}c_{s}^{3}}\int_{0}^{\pi}&|\tilde{\psi}\Big{(}\frac{\omega}{c_{s}}\sin{\theta},\frac{\omega}{c_{s}}\cos{\theta}\Big{)}|^{4}\times\\ &[1-J_{0}\Big{(}\frac{\omega d}{c_{s}}\sin{\theta}\Big{)}]\sin{\theta}d\theta,\end{split}$ (E39)
@@
$\begin{split}F(\omega)=\frac{v\omega^{2}}{\pi^{2}c_{s}^{3}}\int_{0}^{1}&\Big{|}\tilde{\psi}\Big{(}\frac{\omega}{c_{s}}\sqrt{1-x^{2}},\frac{\omega}{c_{s}}x\Big{)}\Big{|}^{4}\times\\ &\Big{[}1-J_{0}\Big{(}\frac{\omega d}{c_{s}}\sqrt{1-x^{2}}\Big{)}\Big{]}dx.\end{split}$ (E40)
Assuming a sharply peaked electron density simplifies the integral as $|\tilde{\psi}(\bm{\kappa},k_{z})|^{4}\approx 1$. In the context of a double quantum dot system this is a reasonable assumption as the dots are usually confined to a narrow region of space in all directions. The finite extent of the dot wavefunctions, assumed to be Gaussian $\psi(\mathbf{r})\propto e^{-r^{2}/2a^{2}}$, introduces a Gaussian decay from the Fourier transform of the dot wavefunction in E37.
@@
$F(\omega)=\frac{v}{\pi^{2}cd^{2}}\bigg{(}\frac{\omega}{\omega_{d}}\bigg{)}^{2}\bigg{(}1-\operatorname{sinc}{\frac{\omega}{\omega_{d}}}\bigg{)}e^{-\omega^{2}/2\omega_{a}^{2}},$ (E41)
@@
$M_{\mathbf{k}}=\bigg{(}\frac{1}{2M\omega_{\mathbf{k}}}\bigg{)}^{-1/2}(D|\mathbf{k}|+iP),$ (E42)